OpenStudy (anonymous):

How many solutions does the equation 6n-6n-12=14 - 2 have ? A. Two B. One C. None D. Infinite

pooja195 (pooja195):

m=m <---infinite solutions m=x <----no solution m=1 <----one solution m= a number x= a number

OpenStudy (anonymous):

So what, infinite ?

pooja195 (pooja195):

Have you solved the problem yet? Or attempted to do it yet?

OpenStudy (anonymous):

I looked at my notes that i have and it says that its NONE.

OpenStudy (anonymous):

6n-6n-12=14-2 -12= 12 this is what i did